Which quadrant contains the vertex of the following: f (x) = 2x^2 – 8x + 11?
Please explain.

Answer by scott8148(6628)

the general vertex form of the quadratic is ___ y = a(x - h)^2 + k ___ this is a parabola with the vertex at (h,k)

y = 2(x^2 - 4x) + 11

completing the square (and compensating so the equation is unchanged) ___ y = 2(x^2 - 4x + 4) - 8 + 11

y = 2(x - 2)^2 + 3